# Blockchain Economic Denial of Sustainability Attack: Exploiting Latency Optimization in Ethereum Transaction Forwarding

### v1

@@ 9.2 Limitations and mitigations
Second, while our method of estimating the number of active peers requires the assumption of the network being static over time, we mitigate this concern by only including the nodes with a significant number of messages—stable nodes, thereby reinforcing the validity of our assumption.
@@
Possible mitigations. Previous work suggests a few countermeasures to prevent EDoS attack [5]: 1) testing the client (Turing test, cryptographic puzzle) to exclude non-human requests or 2) blocklisting users by detecting abnormal traffic. Those solutions do not directly translate to our context because 1) a blockchain node is not interactive (i.e., a human would not be able to answer challenges at the requested rate) and 2) node identity can be easily spoofed. We devise three solutions tailored to blockchain P2P networks.
@@
Second, one could postpone the validation process by introducing a two-step propagation mechanism. The idea is similar to the solutions proposed by Das et al [11] for delaying block validation in the context of the Verifier’s Dilemma. After the modified node receives transactions, it 1) directly relays them to users without validation, ensuring latency reduction, 2) validates transactions afterward, and 3) sends them to the rest (or does not forward them at all). This two-step approach does not incur additional egress traffic. However, modified nodes still pay the cost of delivering invalid transactions to users with an unreliable txpool view, creating a risk for users to include invalid transactions.
@@ 10.2 P2P network security: DoS and spam
Due to the rise of centralized services such as mining pools, exchanges, and relay nodes, Denial of Service (DoS) attacks on those services can have a disastrous impact on the network [49]. The most directly relevant work is by Li et al [35] (and its extension by Yaish et al [54]), which exploits the txpool’s propagation policy by sending transactions with a future nonce to evict legitimate transactions in the txpool. Further, Wang et al [53] developed a fuzzer to automatically discover txpool vulnerabilities. Zhou et al [56] document the spamming behavior of MEV bots as a computational overhead to the network. Our paper describes a new Blockchain DoS attack, exploiting the lack of transaction validations—a superset of the above DoS-related works, as shown in §7), and empirically derives the amplification factors. We also reference some metrics from non-blockchain DoS literature such as amplified DoS and EDoS attacks from Kumar [32] and Wang et al. [52], respectively.
@@ 11 Conclusion
A reduction in latency creates an economic value as a form of MEV/BEV, and some entities modify the software client (“modified node”) to shorten the transaction validation process. We formalize a Blockchain EDoS (Economic Denial of Sustainability) attack, identify similar attack instances in the wild and empirically measure and simulate the attack impact. Our model concludes that an attacker can amplify its original traffic by 3,600 times, and the financial loss by 13,800 times. A lack of transaction validation not only incurs a financial loss on the modified nodes but also causes an uptick in the number of invalid transactions in the network. This degrades the user experience, creates negative externalities by impacting other nodes, and poses a threat to the overall stability of the blockchain ecosystem.
@@ Appendix C Manual analysis of the empirical attack
We manually look at some attackers’ addresses and list two possible motivations for the attack. The first interpretation is to disrupt the service or the network. If the account sends many transactions with insufficient addresses or past nonce, the attacker has no intention of making transactions on-chain, but solely to cause a disturbance to the service/network (i.e., EDoS attack).
@@
Despite different attack intentions (or scale), the attacker crafts many invalid transactions and disrupts the service/network. The prevalence of the attack in the current P2P network corroborates the practicality/feasibility of our EDoS attack.
@@ Appendix E Customized nodes
We find a cluster of nodes that customize the existing clients and collectively develop the software privately. As noted in §LABEL:subsec:other_data, we collect the node name to profile P2P network nodes. Node names (e.g., Geth/v1.13.4-stable-3f907d6a/linux-amd64/go1.21.3) generally follow a convention indicating 1) software name, 2) version/branch and the first 4 bytes of the last git commit hash, 3) operating system, and 4) programming language and its version. We use regular expressions to extract the version and git commit hash. If the node owner forks existing projects, modifies any part of the code, and updates the repository, a different git hash is generated. To determine whether the node is customized, we call the Github API [^16] https://docs.github.com/en/rest/commits/commits?apiVersion=2022-11-28 to check whether the commit publicly exists in the corresponding software repository. We call nodes with non-public git commit hashes customized nodes. Given the space of possible hashes (4 bytes: $16^{8}$), the probability of generating two identical hashes by mere coincidence is extremely low, which suggests the cluster of nodes being operated by the same entity.
@@ Appendix G Attack scenario
We aim to estimate the highest possible traffic costs that modified nodes would incur when an attacker saturates their outgoing traffic with its invalid transactions. The attacker can utilize 400 Ethereum blockchain accounts (as shown in Figure 10) to evict existing transactions in the txpool and keep refreshing it with the new invalid ones, causing the modified nodes to perpetually forward invalid transactions to their neighbors. The attacker can also send transactions from multiple nodes with different IPs, and distribute the attack to multiple modified nodes to scale up the attack until the bandwidth limit (i.e., Distributed EDoS attack). This helps attackers circumvent any network management tools deployed by modified nodes (e.g., anomaly detection—one of the solutions to classical EDoS attacks). We partly reference Kumar’s work [32] to calculate the total link capacity of intermediate nodes (i.e., modified nodes in our case). Table IV summarizes the assumptions and the final cost of modified nodes. We use the bandwidth of AWS EC2 instances for attacker/modified nodes and the minimum bandwidth requirement by Geth for the regular nodes. [^17] https://geth.ethereum.org/docs/getting-started/hardware-requirements In our case, the scale of the attack (the amount of waste) is mainly constrained by the number/bandwidth of modified/regular nodes since the attacker’s traffic gets easily amplified by TAF. Based on our calculation, each modified node could spend at most 88,904 USD, or roughly 8M USD per month in aggregate if all the modified nodes are from the same entity.

### v2

@@ 9.2 Limitations and mitigations
Second, our method of estimating the number of active peers requires the assumption of the network being static over time. We mitigate this concern by only including the nodes with a significant number of messages—stable nodes, thereby reinforcing the validity of our assumption.
@@
Possible mitigations. Previous work suggests a few countermeasures to prevent DoS/EDoS attack [5]: 1) testing the client (Turing test, cryptographic puzzle) to exclude non-human requests or 2) blocklisting users by detecting abnormal traffic. Those solutions do not directly translate to our context because 1) a blockchain node is not interactive (i.e., a human would not be able to answer challenges at the requested rate) and 2) node identity can be easily spoofed. Unlike previous blockchain DoS attacks [36, 55, 28], our identifid vulnerability does not exist in the public client, thereby patching the current client does not resolve the issue. We devise three solutions tailored to our case.
@@
Second, one could postpone the validation process by introducing a two-step propagation mechanism. The idea is similar to the solutions proposed by Das et al [11] for delaying block validation in the context of the Verifier’s Dilemma. After the modified node receives transactions, it 1) directly relays them to users without validation, ensuring latency reduction, 2) validates transactions afterward, and 3) sends them to the rest (or does not forward them at all). While this two-step approach does not negatively impacts the rest of the network, modified nodes still pay the cost of delivering invalid transactions to users, creating a risk for users to include invalid transactions.
@@ 10.2 P2P network security: DoS and spam
Due to the rise of centralized services such as mining pools, exchanges, and relay nodes, Denial of Service (DoS) attacks on those services can have a disastrous impact on the network [50]. The most directly relevant work is by Li et al [36] (and its extension by Yaish et al [55]), which exploits the txpool’s propagation policy by sending transactions with a future nonce to evict legitimate transactions in the txpool. Further, Wang et al [54] developed a fuzzer to automatically discover txpool vulnerabilities. Heo et al [28] illustrate that attackers can delay honest transaction execution by sending invalid transactions. Zhou et al [57] document the spamming behavior of MEV bots as a computational overhead to the network. Our paper describes a new Blockchain DoS attack, exploiting the lack of transaction validations—a superset of the above DoS-related works, as shown in §7), and empirically derives the amplification factors. We also reference some metrics from non-blockchain DoS literature such as amplified DoS and EDoS attacks from Kumar [33] and Wang et al. [53], respectively.
@@ 11 Conclusion
A reduction in latency creates an economic value as a form of MEV/BEV, and some entities modify the software client (“modified node”) to shorten the transaction validation process. We formalize the Blockchain Amplification Attack, identify similar attack instances in the wild and empirically measure and simulate the attack impact. Our model concludes that an attacker can amplify its original traffic by 3,600 times, and the financial loss by 13,800 times. A lack of transaction validation not only increases traffic costs on the modified nodes but also introduce invalid transactions, and degrades the user experience. Our modeling frameworks provide a foundation for explaining how attack transactions propagates through the network and pose a threat to the overall blockchain ecosystem.
12 Ethics
Following Tang et al. [49] on ethics of measuring blockchain P2P networks, our work consists of passive measurements (i.e., not submitting transactions). We do collect the Ethereum node IP addresses, but this is already publicly available information. As a part of responsible disclosure, we contacted the services that we discovered are running modified nodes and could be victims of this attack. One of the services appears to have fixed the vulnerability.
@@ Appendix C Manual analysis of the empirical attack
We manually look at some attackers’ addresses and list two possible motivations for the attack. The first interpretation is to disrupt the service or the network. If the account sends many transactions with insufficient addresses or past nonce, the attacker has no intention of making transactions on-chain, but solely to cause a disturbance to the service/network (i.e., Amplification attack).
@@
Despite different attack intentions (or scale), the attacker crafts many invalid transactions and disrupts the service/network. The prevalence of the attack in the current P2P network corroborates the practicality/feasibility of our proposed attack.
@@ Appendix E Customized nodes
We find a cluster of nodes that customize the existing clients and collectively develop the software privately. As noted in §3.2, we collect the node name to profile P2P network nodes. Node names (e.g., Geth/v1.13.4-stable-3f907d6a/linux-amd64/go1.21.3) generally follow a convention indicating 1) software name, 2) version/branch and the first 4 bytes of the last git commit hash, 3) operating system, and 4) programming language and its version. We use regular expressions to extract the version and git commit hash. If the node owner forks existing projects, modifies any part of the code, and updates the repository, a different git hash is generated. To determine whether the node is customized, we call the Github API [^16] https://docs.github.com/en/rest/commits/commits?apiVersion=2022-11-28 to check whether the commit publicly exists in the corresponding software repository. We call nodes with non-public git commit hashes customized nodes (not modified nodes). Given the space of possible hashes (4 bytes: $16^{8}$), the probability of generating two identical hashes by mere coincidence is extremely low, which suggests the cluster of nodes being operated by the same entity.
@@ Appendix G Attack scenario
We aim to estimate the highest possible traffic costs that modified nodes would incur when an attacker saturates their outgoing traffic with its invalid transactions. The attacker can utilize 400 Ethereum blockchain accounts (as shown in Figure 10) to evict existing transactions in the txpool and keep refreshing it with the new invalid ones, causing the modified nodes to continuously forward invalid transactions to their neighbors. The attacker can also send transactions from multiple nodes with different IPs, and distribute the attack to multiple modified nodes to scale up the attack until the bandwidth limit (i.e., Distributed EDoS attack). This helps attackers circumvent any network management tools deployed by modified nodes (e.g., anomaly detection—one of the solutions to classical EDoS attacks). We partly reference Kumar’s work [33] to calculate the total link capacity of intermediate nodes (i.e., modified nodes in our case). Table IV summarizes the assumptions and the final cost of modified nodes. We use the bandwidth of AWS EC2 instances for attacker/modified nodes and the minimum bandwidth requirement by Geth for the regular nodes. [^17] https://geth.ethereum.org/docs/getting-started/hardware-requirements In our case, the scale of the attack (the amount of waste) is mainly constrained by the number/bandwidth of modified/regular nodes since the attacker’s traffic gets easily amplified by TAF. Based on our calculation, each modified node could spend at most 88,904 USD, or roughly 8M USD per month in aggregate if all the modified nodes are from the same entity.